### v2

# Synthetic POMDPs to Challenge Memory-Augmented RL: Memory Demand Structure Modeling

Abstract
Recent research has developed benchmarks for memory-augmented reinforcement learning (RL) algorithms, providing Partially Observable Markov Decision Process (POMDP) environments where agents depend on past observations to make decisions. While many benchmarks incorporate sufficiently complex real-world problems, they lack controllability over the degree of challenges posed to memory models. In contrast, synthetic environments enable fine-grained manipulation of dynamics, making them critical for detailed and rigorous evaluation of memory-augmented RL. Our study focuses on POMDP synthesis with three key contributions: 1. A theoretical framework for analyzing POMDPs, grounded in Memory Demand Structure (MDS), transition invariance, and related concepts; 2. A methodology leveraging linear process dynamics, state aggregation, and reward redistribution to construct customized POMDPs with predefined properties; 3. Empirically validated series of POMDP environments with increasing difficulty levels, designed based on our theoretical insights. Our work clarifies the challenges of memory-augmented RL in solving POMDPs, provides guidelines for analyzing and designing POMDP environments, and offers empirical support for selecting memory models in RL tasks.
1 Introduction
Within the domain of reinforcement learning (RL), the Partially Observable Markov Decision Process (POMDP) provides a standard formalism for simulating decision-making in practical settings, requiring agents to deduce latent system states from historical data. RL agents enhanced with memory components, for instance Long Short-Term Memory networks (Hochreiter and Schmidhuber 1997), represent a principal strategy for addressing POMDPs. The incorporation of memory allows these agents to integrate historical observations, thus re-establishing guarantees for policy convergence within POMDP contexts (Sutton and Barto 2018).
A variety of benchmarking suites have been developed in prior work to assess RL algorithms across numerous facets (Bellemare et al. 2013; Brockman et al. 2016; Osband et al. 2019; Küttler et al. 2020; Cobbe et al. 2020; Raffin et al. 2021; Hafner 2021; Morad et al. 2023; Huang et al. 2024; Tao et al. 2025). The significant presence of POMDP-based tasks within these benchmarks indicates a growing research focus on addressing POMDPs using RL techniques. These suites incorporate traditional control issues, challenges derived from games, and other settings, a number of which are artificially constructed.
Synthetic environments are crafted for focused evaluation, featuring designs like obscuring or adding noise to true states to test an agent’s ability to handle incomplete information, implementing sparse rewards to gauge exploration strategies, or creating tasks that specifically probe memory recall. Although environments derived from real-world applications possess inherent complexity suitable for testing the overall efficacy of memory-augmented RL, synthetic settings are still crucial. Their straightforward and adjustable nature offers researchers enhanced flexibility to shape task mechanics, free from the entangled intricacies found in real scenarios. This capacity for precise control allows for isolated and directed measurement of particular algorithmic capabilities. Owing to the varied nature of practical challenges and the absence of an agreed-upon standard memory architecture for RL, such focused evaluations maintain substantial importance.
Despite the widespread incorporation of artificial POMDP settings in contemporary benchmarks, significant shortcomings remain. Firstly, the lack of a conceptual structure to define the challenge a POMDP poses to memory models means that creating new environments still depends heavily on heuristic judgment. This casts doubt on whether existing synthetic tasks sufficiently represent a wide spectrum of Memory Demand Structures (MDS) and complexity gradients. Secondly, a systematic procedure for building synthetic POMDPs to test memory-augmented RL is missing, coupled with inadequate precision in tuning environmental attributes.
Presently, two main strategies are predominantly used to vary the difficulty of synthetic POMDPs: 1. Regulating the observable state information through masking techniques (Raffin et al. 2021; Morad et al. 2023; Tao et al. 2025). 2. Altering the problem scope (such as episode length, dimensionality of state/action spaces) or introducing stochasticity (Osband et al. 2019; Cobbe et al. 2020; Morad et al. 2023; Tao et al. 2025). The first strategy, while straightforward to apply, does not adequately reflect the relative importance of different information elements for decision-making. As a result, the difficulty levels produced are often few and not uniformly spaced. The second approach generates multiple difficulty tiers from a base problem, but any performance drop with increased difficulty cannot be uniquely ascribed to memory model shortcomings; limitations of the core RL algorithm might also be a factor. This is because these methods change the fundamental belief MDP (the MDP defined over belief states in a POMDP) (Kaelbling, Littman, and Cassandra 1998) concurrently with altering the MDS. The belief MDP quantifies a problem’s inherent difficulty for an RL agent, separate from the challenge of inferring belief states from trajectories. For a clean evaluation of RL memory models, difficulty adjustments should, ideally, keep the underlying belief MDP invariant. Our examination of the literature found no prior work successfully achieving this.
This research seeks to overcome the stated limitations by introducing a structured framework for generating POMDPs tailored to evaluating memory models in RL. The key contributions of this paper are:
An innovative theoretical framework
for analyzing POMDPs based on MDS and transition invariance (encompassing stationarity and consistency). MDS quantifies the memory requirements dictated by a trajectory, whereas consistency and stationarity describe the fitting demands placed on memory models across observational and temporal scales.
A general construction methodology
that utilizes linear process dynamics, state aggregation, and reward redistribution, allowing for the creation of POMDPs with predefined memory challenges. Herein, linear process dynamics builds POMDPs directly; state aggregation and reward redistribution derive POMDPs from pre-existing MDPs, with state aggregation maintaining the original belief MDP and reward redistribution preserving the optimal policy.
POMDP environment series
that are empirically validated, showing their utility in discriminating the performance of different memory models used in RL.
Grounded in theoretical insights and experimental proof, our proposed method supplies practical guidance for crafting POMDP environments, establishing a solid basis for the assessment of memory models in RL.
2 Related Work
Existing benchmarks offer diverse environments for evaluating RL algorithms. Among them, synthetic POMDPs stand out for their precisely controlled dynamics and adjustable complexity, enabling fine-grained evaluation. These synthetic POMDPs can be systematically classified into two types: those constructed from scratch and those derived from existing MDPs. Our literature review concentrates on these two classes of synthetic environments, providing a concise overview of their key design considerations.
2.1 Environments Constructed from Scratch
Synthetic environments constructed from scratch have a long history in evaluating reinforcement learning algorithms and their use continues to this day. For example, the T-maze, derived from neuroscience experiments, tests the duration of memory for a single bit (O’Keefe and Dostrovsky 1971; Bakker 2001; Osband et al. 2019; Tao et al. 2025), while its variant MemoryLength assesses the number of bits that can be remembered (Osband et al. 2019). Other examples include RockSample, which tests active information collection and coping with object uncertainty (Smith and Simmons 2004; Tao et al. 2025); DeepSea, designed to construct shallow traps for evaluating exploration ability (Osband et al. 2019); Battleship, where hidden information is gradually revealed during gameplay to test reasoning regarding spatial uncertainty (Silver and Veness 2010; Tao et al. 2025); DiscountingChain and Umbrella, which assess credit assignment in trajectories (Osband et al. 2019); and RepeatFirst, RepeatPrevious, Autoencode, CountRecall, and Concentration, which test memory ability through specific recall tasks (Morad et al. 2023).
2.2 Environments Derived from Existing
Constructing new environments from existing ones is a common practice. Researchers typically derive POMDP environments from existing MDP environments by state masking, introducing noise, reward shaping, etc.
Examples of state masking include StatelessCartPole and StatelessPendulum (Morad et al. 2023), which hide position and angle in classic control problems; MaskedMujoco, which retains only velocity observations in Mujoco environments (Todorov, Erez, and Tassa 2012); VisualMujoco (Han, Doya, and Tani 2019; Tao et al. 2025), which replaces joint states with images; and NoInventoryCrafter, which uses image observations and removes agent inventory information in Crafter (Hafner 2021).
Examples of noise introduction include NoisyCartPole and NoisyPendulum, which simulate Gaussian noise in physical environments; and MultiarmedBandit (Morad et al. 2023), which randomizes each episode to assess an agent’s exploration-exploitation trade-off.
Reward shaping, a technique commonly used to accelerate the convergence of the RL algorithm, modifies the reward function to transform it from sparse to dense and from high-variance to smooth, facilitating the learning of optimal or suboptimal policies. Notable conclusions in this regard include that potential-based reward shaping does not alter the set of optimal policies (Ng, Harada, and Russell 1999), among others. From an environmental perspective, this technique serves to adjust the difficulty of the environment.
3 Preliminaries
MDP is the standard theoretical model in RL, where decisions depend solely on the current state. By contrast, POMDP models scenarios with hidden information, such that decisions may rely on all past observations.
Definition 3.1 (Markov Decision Process (MDP)).
An MDP $\mathcal{M}\coloneqq\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ consists of: initial state distribution $\rho_{0}\in\Delta_{S}$; state set $S$; action set $A$; transition and reward dynamics at time $t$: $T_{t}:S\times A\to\Delta_{S\times\real}$.
Definition 3.2 (Partially Observable MDP (POMDP)).
A POMDP $\mathcal{N}\coloneqq\langle\rho_{0},S,A,Z,\{T_{t}\}_{t=0}^{\infty},\{O_{t}\}_{t=0}^{\infty}\rangle$ consists of: initial state distribution $\rho_{0}\in\Delta_{S}$; state set $S$; action set $A$; observation set $Z$; transition and reward dynamics at time $t$: $T_{t}:S\times A\to\Delta_{S\times\real}$; observation function: $O_{t}:S\to\Delta_{Z}$.
Unless otherwise stated, in subsequent discussions, we assume $S$ and $Z$ are compact subsets of k. For ease of subsequent analysis, we introduce HDP, which omits the state concept and defines transitions using observations.
Definition 3.3 (History Dependent Decision Process (HDP)).
An HDP $\mathcal{H}\coloneqq\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ consists of: initial observation distribution $\rho_{0}\in\Delta_{Z}$; observation set $Z$; action set $A$; transition and reward dynamics at time $t$: $T_{t}:Z^{t+1}\times A^{t+1}\to\Delta_{Z\times\real}$.
In MDPs, a policy maps a state to an action via $\pi:S\to\Delta_{A}$. In contrast, POMDP and HDP policy maps past observations and actions to an action via $\pi:(Z\times A)^{\ast}\to\Delta_{A}$. Compared with POMDP, which models via observation functions and states, HDP directly models the relationship between the next observation-reward pair and all previous observations and actions, granting us greater freedom to construct transition mechanisms. However, can HDP accurately represent the transition dynamics of POMDP? We provide a positive answer to this question with the following theorem. The proof is provided in Appendix A.
Theorem 3.1 (POMDP’s Equivalent HDP).
For any POMDP $\mathcal{N}=\langle\rho_{0},S,A,Z,\{T_{t}\}_{t=0}^{\infty},\{O_{t}\}_{t=0}^{\infty}\rangle$, it has equivalent HDP $\mathcal{H}=\langle\rho_{0}^{\prime},Z^{\prime},A^{\prime},\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$ where $Z^{\prime}=Z$, $A^{\prime}=A$, $\rho_{0}^{\prime}(z)=\sum_{s\in S}\rho_{0}(s)O_{0}(z|s)$, and $T_{t}^{\prime}(z_{t+1},r_{t}|z_{0:t},a_{0:t})=$
$\frac{\sum\limits_{s_{0:t+1}\in S^{t+2}}\!\!\!\!\!\!\!\!\!T_{t}(s_{t+1},r_{t}|s_{t},a_{t})O_{t+1}(z_{t+1}|s_{t+1})\Phi_{t}(z_{0:t}|s_{0:t};a_{0:t})}{\sum\limits_{s_{0:t}\in S^{t+1}}\!\!\!\!\!\!\!\Phi_{t}(z_{0:t}|s_{0:t},a_{0:t})}$
where $\Phi_{t}(z_{0:t}|s_{0:t},a_{0:t})=$
$\rho_{0}(s_{0})O_{0}(z_{0}|s_{0})\!\!\prod_{\tau=0}^{t-1}\!O_{\tau+1}(z_{\tau+1}|s_{\tau+1})\!\!\!\sum_{r_{\tau}\in\real}\!\!T_{\tau}(s_{\tau+1},\!r_{\tau}|s_{\tau},\!a_{\tau})$
The concept of equivalence in the Thm. 3.1 describes the fact that for any given POMDP, there exists a corresponding HDP such that, from the perspective of any agent with no knowledge of the dynamics of the decision process, the HDP is indistinguishable from the POMDP.
4 Memory Demand Structure (MDS)
In solving POMDPs, belief states are often used instead of the entire trajectory as the basis for an agent’s decisions. This is because, like the entire trajectory, a belief state qualifies as a sufficient statistic (Kaelbling, Littman, and Cassandra 1998): no additional past actions or observations can provide further information about the current world state (White III and Scherer 1989). A sufficient statistic serves as the basis for decision-making and requires the agent to retain relevant information from the trajectory. However, not all state-action pairs at every time step are useful. There are ”key events” that can completely determine the next transition and reward. Building on this insight, we introduce the concept of MDS, which refers to the critical time steps within a trajectory that the agent has to memorize to predict the next state and reward.
Definition 4.1 (Memory Demand Structure (MDS)).
For any trajectory $h_{t}=(z_{0:t},a_{0:t},r_{0:t-1})$ of a POMDP $\mathcal{N}=\langle\rho_{0},S,A,Z,\{T_{t}\}_{t=0}^{\infty},\{O_{t}\}_{t=0}^{\infty}\rangle$, a set $D\subseteq\{0,1,\cdots,t\}$ is called $h_{t}$’s MDS if $\forall s_{t+1}\in S,~\forall r_{t}\in\real$,
$P(s_{t+1},r_{t}|z_{0:t},a_{0:t})\!=\!P(s_{t+1},r_{t}|\bigcap_{\tau\in D}\{Z_{\tau}\!=\!z_{\tau},A_{\tau}\!=\!a_{\tau}\})$
We denote $\mathbf{D}(h_{t})$, the set of all MDSs of $h_{t}$; and $\tilde{D}(h_{t})\coloneqq\bigcap_{\tau\in D}\{Z_{\tau}=z_{\tau},A_{\tau}=a_{\tau}\}$, the event sufficient for estimating next state $s_{t+1}$ and reward $r_{t}$ joint distribution. MDSs can also be defined for HDP. Since HDP has no explicit state concept, both sides of the equation in Def. 4.1 are replaced with the conditional joint distribution of next observation and reward, i.e., $P(z_{t+1},r_{t}|z_{0:t},a_{0:t})=P(z_{t+1},r_{t}|\tilde{D}(h_{t}))$. A detailed discussion of MDSs for POMDPs and their HDP equivalents is provided in Appendix B.1-B.2.
It is worth noting that for the same trajectory of a POMDP, there may be different MDSs. These structures are not necessarily inclusive of each other and may even be completely disjoint. Appendix B.3 presents such a POMDP.
For any trajectory $h_{t}$, $\{0,1,\dots,t\}\in\mathbf{D}(h_{t})$ always holds, but memorizing all state-action pairs at each timestep is not a universal requirement of POMDPs for memory models. For the same trajectory $h_{t}$ with multiple MDSs, when analyzing the required capabilities of an RL memory model, we should select appropriately from $\mathbf{D}(h_{t})$, prioritizing ”simpler” ones, such as those with the minimum number of timesteps (denoted $D^{M}$) or where the oldest timestep is closest to the current time $t$ (denoted $D^{C}$). These criteria derive from the characteristics of major memory models: limited capacity and a tendency to remember recent information while forgetting older details. Furthermore, when using these criteria to select MDSs, trajectories of equal length within the same POMDP may share the same MDS or differ, depending on the POMDP’s properties.
Figure 1: Visualization of MDSs of some example synthetic POMDP environments in Section 2. A colored block $(t,i)$ indicates that for any trajectory $h_{t}$, $i\in D^{M}\in\mathbf{D}(h_{t})$. Different colors denote distinct dependence relationships: state-only, action-only, and both. Details in Appendix B.4.
In the POMDPs illustrated in Fig. 1, all trajectories of equal length share identical MDSs. Subsequent sections focus on synthetic environments with this uniformity, as it enables a more intuitive demonstration of different memory models’ capability spectra. However, uniform MDSs do not mean that probability distributions estimated from equal-length trajectories depend identically on past timesteps within the MDS. The next section details how MDS relates to designing specific environmental transition dynamics.
5 Constructed from Scratch: MDS, Invariance, and Dynamics
This section introduces our proposed method for constructing HDP environments from scratch, which involves three steps: 1. selecting appropriate MDSs; 2. determining transition invariance, including stationarity and consistency; 3. specifying the exact transition and reward dynamics.
5.1 Selecting MDS
When constructing an HDP environment, a natural question arises: what constitutes its difficulty, and how can such difficulties be instantiated? For MDP environments, this question admits multiple answers, including randomness, large state/action spaces, and sparse reward distributions. However, even with these various sources of complexity, the MDS for any trajectory $h_{t}$ of an MDP remains extremely simple: $\{t\}$. In contrast, HDPs inherently possess complex MDS structures absent in MDPs. Since HDPs retain all the aforementioned difficulty factors, constructing diverse MDS forms should be the primary focus when building HDP environments. To simplify subsequent discussions, we assume that trajectories of the same length share identical MDSs, and MDSs of trajectories with different lengths are translation-invariant. Specifically, we designed a series of high-order MDPs (see Appendix C.3) whose transition and reward dynamics depend on both the current state-action pair and states from several preceding moments (Fig. 2). Increasing the number of preceding moments in the MDS makes the constructed environment more challenging.
Figure 2: Visualization of MDSs of selected high-order MDPs, with order increasing from 1 to 4. Only the current action and the most recent several states affect transitions.
5.2 Determining Invariance
While MDS identifies the moments influencing predictions of the joint distribution of next-step observations and rewards, it fails to capture how these moments exert their influence. The operational manners of these moments can be highly complex and vary across trajectories, making this our second focus. To better constrain the difficulty of HDPs, we need to restrict the diversity of such operational manners.
We note that in Independent Decision Processes (IDP; see Appendix C.1), such as the multi-armed bandit problem, the joint distribution of next observation and reward is identical across all trajectories. In stationary MDPs (see Appendix C.2), this distribution is the same for trajectories with identical final observations and actions. In non-stationary MDPs, it is identical for trajectories of the same length that share identical final observations and actions. Extending this line of thinking, we arrive at the following concept.
Definition 5.1 (HDP’s Transition Invariance).
Given an HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ and an equivalence relation $\simeq$ on its trajectory set $H$, $\mathcal{H}$ is called $\simeq$-invariant if for any $h_{t_{1}},h^{\prime}_{t_{2}}\in H$ such that $h_{t_{1}}\simeq h^{\prime}_{t_{2}}$, the following holds: $\forall z\in Z,~\forall r\in\real,~T_{t_{1}}(z,r|z_{0:t_{1}},a_{0:t_{1}})=T_{t_{2}}(z,r|z_{0:t_{2}}^{\prime},a_{0:t_{2}}^{\prime})$.
All equivalence relations on $H$ form a lattice (see Appendix C.5). We denote by $\simeq^{\bot}$ the finest equivalence on $H$, where distinct trajectories are inequivalent. The universal equivalence $\simeq^{\top}$ treats all trajectories as equivalent. We use $\simeq^{n}$ for the relation where trajectories $h_{t_{1}},h_{t_{2}}^{\prime}$ are equivalent if $\lfloor t_{1}/n\rfloor=\lfloor t_{2}/n\rfloor$, $\simeq_{k}$ for the relation where trajectories are equivalent if their most recent $k$ state-action pairs are identical. With definitions above, IDPs are $\simeq^{\top}$-invariant HDPs, stationary MDPs are $\simeq_{1}$-invariant, non-stationary MDPs are $(\simeq_{1}\wedge\simeq^{1})$-invariant, general HDPs are $\simeq^{\bot}$-invariant. In general, finer equivalence relations induce more diverse joint distributions of next-step observations and rewards, making the HDP more challenging.
We use the term (intra-trajectory) stationarity to describe the invariance of the transition dynamics across steps within the same trajectory, and (inter-trajectory) consistency to describe the invariance of the transition dynamics among different trajectories (see Appendix C.6). Using the above notations, we derive four types of HDPs as presented in Tab. 1, with visualizations in Fig. 3 (see Appendix C.7).
Invariance Stationary Non-stationary
Consistent $\simeq_{k}$ $\simeq_{k}\wedge\simeq^{n}$
Non-consistent $\simeq_{k}\wedge\simeq^{\lfloor mz_{0}\rfloor}$ $\simeq_{k}\wedge\simeq^{n}\wedge\simeq^{\lfloor mz_{0}\rfloor}$
Table 1: A construction of HDPs ($S=[0,1)$) with different transition invariance properties. $k,n,m\in\mathbb{N}^{+}$, $\simeq^{\lfloor mz_{0}\rfloor}$ denotes an equivalence relation where trajectories $h,h^{\prime}$ are equivalent if their initial observations $z_{0},z_{0}^{\prime}$ lie within the same interval $\left[\frac{i}{m},\frac{i+1}{m}\right)$, $k$ is the order as shown in Fig. 2.
Figure 3: Corresponding to Tab. 1, the x-axis shows $n$-length time periods in trajectories, and the y-axis shows distinct trajectories differentiated by the $\frac{1}{m}$-length intervals that their first observations belong to. Different colors correspond to different transition and reward dynamics.
5.3 Specifying Dynamics
After determining the MDS and transition invariance, details such as whether the next observation emerges as a linear combination of previous ones remain unspecified. Thus, we need to select specific transition dynamics.
We employ a deterministic autoregressive (AR) process (Hamilton 2020), (Fig. 4), with modulo operation as the transition dynamics of the HDP, with initial observations randomly sampled and coefficients $w_{i}^{h_{t}}\geq 1$.
Figure 4: An example AR process with coefficients $w_{0},w_{1}$. $z_{0},~z_{1}$ is sampled independently from $U(0,1)$.
$z_{t}\sim U(0,1),~(t<k)$
$z_{t+1}=\left(\sum_{i=0}^{k-1}w_{i}^{h_{t}}z_{t-i}\right)\mod 1,~(t+1\geq k)$
Agents receive rewards for predicting the next observation to fall within the correct interval, i.e., $r_{t}=\mathbf{1}_{\{a_{t}=\lfloor mz_{t+1}\rfloor\}}$. The observation space is $Z=\left[0,1\right)$, action space is $A=\{0,1,\dots m-1\}$, where choosing action $i$ corresponds to predicting interval $\left[\frac{i}{m},\frac{i+1}{m}\right)$. In non-stationary scenarios, autoregressive coefficients $w^{h_{t}}_{0:k-1}$ vary with $t$ within a single trajectory; in non-consistent scenarios, they vary across different trajectories, where the distinction arises from the intervals in which the initial observations of the trajectories fall. Our design of employing the AR process with modulo operation is motivated by the following considerations:
  • Simplicity: It allows for straightforward implementation with low computational overhead.
  • Scalability: It seamlessly scales difficulty by increasing the order $k$, the action space size $m$, and the observation space dimension (with coefficients extended to matrices).
  • Expressiveness: Linear difference equations effectively model the discretization of numerous physical systems, and their mechanism of generating new observations by aggregating past ones is compatible with any MDS.
  • Boundedness: The modulo operation ensures a compact observation space, ideal for evaluating memory models.
Further details are provided in Appendix D.
6 Derived from Existing: State Aggregation and Reward Redistribution
This section investigates two approaches for constructing HDP environments leveraging existing MDPs. The first aggregates state sequences from MDP trajectories to serve as observations in HDPs. The second adjusts the reward sequences in MDP trajectories to create HDPs. The advantage of our approaches is primarily as follows:
  • Simplicity: Developing wrappers is easier than constructing new HDPs from scratch.
  • Reusability: A single wrapper works across diverse MDP environments.
  • Generality: Existing MDP environments can be reused to construct HDPs, introducing MDS while retaining their inherent characteristics.
  • Interpretability: The methods have clear physical meaning and high interpretability.
Further details are provided in Appendix G.
6.1 State Aggregation
We apply the reversible History Aggregator for State (HAS) (Wang and Li 2025), (Fig. 5), to MDPs to construct HDPs. Specifically, for any trajectory $h_{t}=(s_{0:t},a_{0:t},r_{0:t-1})$ of a given MDP, we construct the HDP observation $z_{t}$ by convolving a fixed sequence $w_{0:t}$ with the MDP states $s_{0:t}$.
$z_{t}\coloneqq w_{0:t}\ast s_{0:t}=\sum_{i=0}^{t}w_{t-i}s_{i}$
Figure 5: A convolution-based HAS (with coefficients $w_{0},w_{1}$) for aggregating MDP states into HDP observations.
Figure 6: Visualization of MDSs of convolution. Left figure coefficients: $1,w$; Right figure coefficients: 1,w1,-w, $(w>0)$. In $W^{-1}_{t}$, positive elements are red, negative blue, and darker colors indicate larger absolute values of the coefficient.
Unlike the dynamics in the previous section, where new observations are generated from prior observations based on an AR process, the convolution-based HAS resembles a moving average (MA) process (Hamilton 2020). However, what it smooths is not a white noise sequence, but a state sequence from a given MDP. When $w_{0}\neq 0$, each $s_{t}$ can be perfectly recovered using the observation sequence $z_{0:t}$ (through deconvolution), this is because $z_{0:t}=W_{t}s_{0:t}$, where $W_{t}$ is an invertible lower triangular Toeplitz matrix. The MDS of each trajectory is determined by the non-zero diagonals of $W_{t}^{-1}$, since $s_{0:t}=W_{t}^{-1}z_{0:t}$. Thus, we can determine the MDS by controlling $W_{t}^{-1}$.
The convolution-based transformation also applies to MDPs with discrete state spaces. For such MDPs (e.g., $S=\{0,1,\dots,N-1\}$), the coefficient $w_{0}$ must be coprime with $N$, and the summation and multiplication operations replaced by their modulo-$N$ counterparts.
This convolution-based transformation has a clear physical meaning: it converts an original echo-free system (MDP) into an HDP with echoes from previous states, which must be removed to restore the original echo-free state. Such a phenomenon is common in signal processing. The belief MDP of the constructed HDP remains identical to the original MDP. See Appendix E for further details.
6.2 Reward Redistribution
The preceding sections of this paper focus primarily on constructing non-Markov state transition mechanisms, while non-Markov rewards also hold considerable significance in related research (Gaon and Brafman 2020; Rens and Raskin 2020; Derman, Dalal, and Mannor 2021; Abel et al. 2021, 2022; Early et al. 2022; Pitis, Bailey, and Ba 2022; Han et al. 2022; Hyde and Santos Jr 2024). In this section, we construct HDPs with non-Markov rewards by modifying the rewards of MDP trajectories. To ensure that such modifications do not violate causality and thus can be implemented as an environmental wrapper, rewards appearing at subsequent moments should not affect rewards from the past. When we merely reallocate rewards in MDP trajectories (e.g., delaying the occurrence of rewards and compensating for time value using discount factors), the optimal policy of the original MDP remains the optimal policy of the HDP (Arjona-Medina et al. 2019) despite the introduction of MDS. This draws on the concept of equivalent cash flows in finance (Jordan, Ross, and Westerfield 2003).
Specifically, we developed a series of wrappers that delay rewards for $k$ steps. For any MDP with discount factor $\gamma$, and trajectory $h_{t}=(s_{0:t},a_{0:t},r_{0:t-1})$ that is not a prefix of any other trajectories, the delayed reward $r^{\prime}_{0:t}$ satisfies:
$r^{\prime}_{0:k-1}=0,~r^{\prime}_{k:t-2}=\frac{r_{0:t-k-2}}{\gamma^{k}},~r^{\prime}_{t-1}=\sum_{i=0}^{k-1}\frac{r_{t-1-i}}{\gamma^{i}}$
Figure 7: Visualization of MDSs of different reward delays. When rewards are overall delayed, the next state still depends on the current state and action, while the reward relies on states and actions from $k$ steps prior. The last reward is the discounted sum of the last $k$ rewards.
The effect of reward delay on MDS is shown in Fig. 7. For a detailed discussion, refer to Appendix F.
7 Experiments
Our experiments were conducted on three sets of environments corresponding to the previous sections: Section 5.3 LinProcEnv series, Section 6.1 StateConv series, and Section 6.2 RewardDelay series. Due to computational constraints, we tested four sequence models: Elman (Elman 1990), LSTM (Hochreiter and Schmidhuber 1997), LMU (Voelker, Kajić, and Eliasmith 2019), and LinearAttention (Katharopoulos et al. 2020), along with one memory-free baseline model, MLP, for the PPO (Schulman et al. 2017) algorithm. All implementations are based on the baselines provided in POPGym (Morad et al. 2023). Each experiment was repeated $10$ times to reduce variance. Combined with the episode average reward curves during training, the total training steps for all experimental groups were unified at $10^{6}$ to ensure convergence. Details are provided in Appendix H.
7.1 LinProcEnv Series
AllEqOneLinProcEnv(k)(k) is a $k$-th order AR process with modulo $1$ operation, where all coefficients equal $1$. {All,Time,Traj,No}EqLinProcEnv(k)(k) is $k$-th order AR process with different coefficients and different transition invariances. MDSs are illustrated in Fig. 2, and consistency/stationarity in Fig. 3 (m=8,n=8m=8,~n=8). These environments have a fixed trajectory length of $64$. At each step, an agent predicts which interval the next observation will fall into, scoring $1$ for correct predictions and 0 otherwise. The maximum reward is thus 64, while random guesses yield a reward close to $64=8$.
Figure 8: Linear processes with coefficients all equal to $1$.
(a) Consistent, Stationary
(b) Consistent, Non-stationary
(c) Non-consistent, Stationary
(d) Non-consistent, Non-stationary
Figure 9: Linear processes with different coefficients and different transition invariance.
Fig. 8 shows the performance degradation of memory models as order increases, under AR processes with all $1$ coefficients. Elman performs significantly better at orders 141–4 but approaches random guesses from order $5$ onward, while the other models near random guesses after order $4$.
Fig. 9 shows the performance under different consistency/stationarity conditions. Elman remains the best-performing model. Notably, under non-consistency conditions, all memory models exhibit significantly poorer performance at every order, whereas non-stationarity has no significant impact on performance. This implies that during PPO training, memory models face greater difficulty in fitting varying transition dynamics among trajectories, whereas temporal non-stationarity has no substantial impact.
7.2 StateConv Series
[…]
RewardWhenInside is an IDP, where states are sampled from $U(0,1)$ for $256$ steps. Agents receive a reward $1$ for correctly identifying the interval (out of $8$) that the current state falls into, otherwise 0. Random guesses yield an episode reward of $256/8=32$. NormalizedCartPole limits the trajectory length of the CartPole to $256$ and normalizes the episode reward to [1,1][-1,1].
(a) RewardWhenInside_p
(b) RewardWhenInside_n
(c) NormalizedCartPole_p
(d) NormalizedCartPole_n
Figure 10: StateConv series with increasing coefficient $w_{1}$.
[…]
7.3 RewardDelay Series
RewardDelay(k)(k) is a series of wrappers that delay rewards for $8k$ steps to construct HDP, with MDSs shown in Fig. 7. Environments with suffix ”_kk” in names are wrapped with it.
Figure 11: RewardWhenInside with increasing delay steps.
Fig. 11 illustrates performance degradation as delay steps increase. Since delayed rewards do not render an MDP’s optimal (Markov) policy non-Markovian, memory models do not outperform the memoryless MLP. LMU and LSTM exhibit the steepest performance drops with more delays, while other models perform on par with the MLP, indicating they are more susceptible to interference from irrelevant observations that should be discarded. Thus, the RewardDelay series evaluates memory models’ ability to forget irrelevant information, perfectly complementing the StateConv series.
8 Conclusion and Future Work
We have established a theoretical framework centered on Memory Demand Structures (MDS) to systematically analyze and generate Partially Observable Markov Decision Processes (POMDPs) with scalable difficulty levels. By integrating the concept of transition invariance—encompassing stationarity and consistency—with methodological tools such as linear process dynamics, state aggregation, and reward redistribution, we constructed three distinct series of POMDP environments. Each series exhibits a progressively increasing level of complexity, providing a structured testbed for evaluating reinforcement learning (RL) agents equipped with diverse memory architectures.
Experimental results revealed a notable trend: the relatively simple Elman network (RNN) frequently outperformed more sophisticated memory models across a majority of the evaluated settings. Furthermore, as the MDS complexity escalated within the environment series, the performance degradation observed in the Elman model was significantly less pronounced compared to other models. This suggests that the Elman network possesses a superior capability in capturing and leveraging higher-order temporal dependencies present in the observation sequences. These findings offer a plausible explanation for the performance advantage of Elman-style networks previously noted in studies such as the POPGym benchmark (Morad et al. 2023).
Based on these insights, we advocate for the inclusion of progressively challenging MDS-based test series as a standard component in the evaluation protocol for memory-augmented RL algorithms. Visualizing performance through curves, such as those depicted in Fig. 8, offers an intuitive means to assess and compare the ability of different memory models to represent complex, long-range dependencies critical for solving RL tasks under partial observability.
Looking forward, our future work will focus on two main directions: 1. Developing a more extensive and varied collection of environment series. This includes designing scenarios where different trajectories within the same environment exhibit heterogeneous MDS patterns, thereby presenting a more nuanced and realistic challenge. 2. Undertaking a broader set of empirical studies aimed at providing practical guidance for model selection. The goal is to establish clearer correlations between specific task characteristics, encoded by MDS profiles, and the most suitable memory model from an expanded set of candidates, ultimately aiding practitioners in optimizing their RL agent designs.
Appendix A Proof of Theorem 3.1: POMDP’s Equivalent HDP
Theorem (3.1).
For any POMDP $\mathcal{N}=\langle\rho_{0},S,A,Z,\{T_{t}\}_{t=0}^{\infty},\{O_{t}\}_{t=0}^{\infty}\rangle$, it has equivalent HDP $\mathcal{H}=\langle\rho_{0}^{\prime},Z^{\prime},A^{\prime},\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$ that: $Z^{\prime}=Z$, $A^{\prime}=A$, $\rho_{0}^{\prime}(z)=\sum_{s\in S}\rho_{0}(s)O_{0}(z|s)$, and
$T_{t}^{\prime}(z_{t+1},r_{t}|z_{0:t},a_{0:t})=\frac{\sum\limits_{s_{0:t+1}\in S^{t+2}}T_{t}(s_{t+1},r_{t}|s_{t},a_{t})O_{t+1}(z_{t+1}|s_{t+1})\Phi_{t}(z_{0:t}|s_{0:t},a_{0:t})}{\sum\limits_{s_{0:t}\in S^{t+1}}\Phi_{t}(z_{0:t}|s_{0:t},a_{0:t})}$
where
$\Phi_{t}(z_{0:t}|s_{0:t},a_{0:t})=\rho_{0}(s_{0})O_{0}(z_{0}|s_{0})\prod_{\tau=0}^{t-1}O_{\tau+1}(z_{\tau+1}|s_{\tau+1})\sum_{r_{\tau}\in\real}T_{\tau}(s_{\tau+1},r_{\tau}|s_{\tau},a_{\tau})$
Proof.
$T_{t}^{\prime}(z_{t+1},r_{t}|z_{0:t},a_{0:t})$
$=\frac{P(Z_{0:t+1}=z_{0:t+1},R_{t}=r_{t}|A_{0:t}=a_{0:t})}{P(Z_{0:t}=z_{0:t}|A_{0:t}=a_{0:t})}$
$=\frac{\sum\limits_{s_{0:t+1}\in S^{t+2}}P(s_{0})P(z_{0}|s_{0})P(s_{t+1},r_{t}|s_{t},a_{t})P(z_{t+1}|s_{t+1})\prod_{\tau=0}^{t-1}P(s_{\tau+1}|s_{\tau},a_{\tau})P(z_{\tau}|s_{\tau})}{\sum\limits_{s_{0:t}\in S^{t+1}}P(s_{0})P(z_{0}|s_{0})\prod_{\tau=0}^{t-1}P({s_{\tau+1}|s_{\tau},a_{\tau})P(z_{\tau}|s_{\tau})}}$
$=\frac{\sum\limits_{s_{0:t+1}\in S^{t+2}}\rho_{0}(s_{0})O_{0}(z_{0}|s_{0})T_{t}(s_{t+1},r_{t}|s_{t},a_{t})O_{t+1}(z_{t+1}|s_{t+1})\prod_{\tau=0}^{t-1}O_{\tau+1}(z_{\tau+1}|s_{\tau+1})\sum\limits_{r_{\tau}\in\real}T_{\tau}(s_{\tau+1},r_{\tau}|s_{\tau},a_{\tau})}{\sum\limits_{s_{0:t}\in S^{t+1}}\rho_{0}(s_{0})O_{0}(z_{0}|s_{0})\prod_{\tau=0}^{t-1}O_{\tau+1}(z_{\tau+1}|s_{\tau+1})\sum\limits_{r_{\tau}\in\real}T_{\tau}(s_{\tau+1},r_{\tau}|s_{\tau},a_{\tau})}$
$=\frac{\sum\limits_{s_{0:t+1}\in S^{t+2}}T_{t}(s_{t+1},r_{t}|s_{t},a_{t})O_{t+1}(z_{t+1}|s_{t+1})\Phi_{t}(z_{0:t}|s_{0:t},a_{0:t})}{\sum\limits_{s_{0:t}\in S^{t+1}}\Phi_{t}(z_{0:t}|s_{0:t},a_{0:t})}$
Appendix B Further Discussion of Section 1. Memory Demand Structure (MDS)
B.1 MDS: Definitions and Extended Connotation
Definition (4.1 POMDP Memory Demand Structure (MDS)).
For any trajectory $h_{t}=(z_{0:t},a_{0:t},r_{0:t-1})$ of a POMDP $\mathcal{N}=\langle\rho_{0},S,A,Z,\{T_{t}\}_{t=0}^{\infty},\{O_{t}\}_{t=0}^{\infty}\rangle$, a set $D\subseteq\{0,1,\cdots,t\}$ is called $h_{t}$’s MDS if $\forall s_{t+1}\in S,~\forall r_{t}\in\real$,
$P(s_{t+1},r_{t}|z_{0:t},a_{0:t})\\ =P(s_{t+1},r_{t}|\bigcap_{\tau\in D}\{Z_{\tau}=z_{\tau},A_{\tau}=a_{\tau}\})$
We denote $\mathbf{D}(h_{t})$, the set of all MDSs of $h_{t}$; $D(h_{t})\coloneqq\{(z_{\tau},a_{\tau})\}_{\tau\in D}$, which is a sufficient statistic; and $\tilde{D}(h)\coloneqq\bigcap_{\tau\in D}\{Z_{\tau}=z_{\tau},A_{\tau}=a_{\tau}\}$, a sufficient condition for estimating current state $s_{t}$.
Since HDP has no explicit concept of hidden Markov states, both sides of the equation in Def. 4.1 should be replaced with the joint distribution of the next observation and reward, which leads to the following definition.
Definition B.1 (HDP Memory Demand Structure (MDS)).
For any trajectory $h_{t}=(z_{0:t},a_{0:t},r_{0:t-1})$ of an HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$, a set $D\subseteq\{0,1,\dots,t\}$ is called $h_{t}$’s MDS if $\forall z_{t+1}\in Z,r_{t}\in\real$,
$T_{t}(z_{t+1},r_{t}|z_{0:t},a_{0:t})=P(z_{t+1},r_{t}|\bigcap_{\tau\in D}\{Z_{\tau}=z_{\tau},A_{\tau}=a_{\tau}\})$
Note: The results below are self-evident.
  • MDP is a special case of POMDP, where $Z=S,\ O_{t}(Z_{t}=s_{t}|S_{t}=s_{t})=1$.
  • For any trajectory $h_{t}$ of a POMDP, we have $\{0,1,\dots,t\}\in\mathbf{D}(h_{t})$.
  • For any trajectory $h_{t}$ of an MDP, we have $\{t\}\in\mathbf{D}(h_{t})$.
  • For any $D\in\mathbf{D}(h_{t})$ and $X\subseteq\{0,1,\dots,t\}$, we have $D\cup X\in\mathbf{D}(h_{t})$.
More Refined Characterization of MDS
For a trajectory $h_{t}$ of an HDP, $D(h_{t})$ is defined to capture the relationship between the next observation-reward pair and past observation-action pairs. This relationship can be decomposed into four components: the relation between the next observation and past observations, between the next observation and past actions, between the next reward and past observations, and between the next reward and past actions. This decomposition can be achieved by taking the marginal distributions of $z_{t+1}$ and $r_{t}$ from $T_{t}(z_{t+1},r_{t}|z_{0:t},a_{0:t})$ along with the technique of separating the observation and action components in $\tilde{D}(h_{t})$.
Such a discussion enables a more refined characterization of MDS.
When we focus on the associations between the next-step state-reward pair (for POMDPs), next-step observation-reward pair (for HDPs), and previous observations within the MDS, our emphasis lies in the memory of the past by the memory model, specifically, which observations provided by the environment have served as clues in the process of reaching the current status. Conversely, when our focus shifts to the associations with previous actions within the MDS, our emphasis turns to the exploration behavior of the RL algorithm: that is, what sequence of actions is required to reach the current status.
In the other dimension, we can define Markov/non-Markov state transition dynamics when focusing solely on the relationship between next state/observation and past observation-action pairs. Conversely, we can also define Markov/non-Markov rewards when concentrating only on the relationship between the next reward and prior observation-action pairs. Consequently, POMDP/HDP can be categorized into four types: Markov-transition-Markov-reward, Markov-transition-non-Markov-reward, Non-Markov-transition-Markov-reward, and Non-Markov-transition-non-Markov-reward.
B.2 MDS: POMDP vs. Equivalent HDP
The MDS of POMDPs is defined using an equation of the conditional joint distribution of the next state and reward, while that of HDPs is defined using an equation of the conditional joint distribution of the next observation and reward. This is natural and convenient: HDPs lack an explicit concept of a current hidden state, making it a necessary substitute to model the next observation-reward joint distribution.
However, this naturally raises a new question: for any trajectory $h_{t}$ shared by a POMDP and an equivalent HDP (as per Theorem 3.1), how do their MDSs differ under the two definitions? We show via the following theorem that although the two MDSs are not necessarily identical, the MDSs of the POMDP fall within those of the HDP.
This conclusion indicates that under such definitions, the MDS of POMDPs tends to encompass more timesteps than that of HDPs, the concept of MDS for HDPs is ”weaker” than that of POMDPs. This aligns with intuition: while the MDS of POMDPs requires predicting the next Markov state and reward, HDPs only need to predict the next observation and reward.
Theorem B.1 (MDS of POMDP and Equivalent HDP).
For any POMDP $\mathcal{N}=\langle\rho_{0},S,A,Z,\{T_{t}\}_{t=0}^{\infty},\{O_{t}\}_{t=0}^{\infty}\rangle$ with its equivalent HDP $\mathcal{H}=\langle\rho_{0}^{\prime},Z,A,\{T^{\prime}_{t}\}_{t=0}^{\infty}\rangle$, if $h_{t}=(z_{0:t},a_{0:t},r_{0:t-1})$ is a trajectory of $\mathcal{N}$ and $\mathcal{H}$, then $\mathbf{D}_{\mathcal{N}}(h_{t})\subseteq\mathbf{D}_{\mathcal{H}}(h_{t})$.
Proof.
For $D_{\mathcal{N}}\in\mathbf{D}_{\mathcal{N}}(h_{t})$, by the definition of POMDP MDS, we have
$P(s_{t+1},r_{t}|z_{0:t},a_{0:t})\\ =P(s_{t+1},r_{t}|\bigcap_{\tau\in D_{\mathcal{N}}}\{Z_{\tau}=z_{\tau},A_{\tau}=a_{\tau}\})$
Then, by the definition of $T^{\prime}_{t}$, we have
$T_{t}^{\prime}(z_{t+1},r_{t}|z_{0:t},a_{0:t})$
$=P(Z_{t+1}=z_{t+1},R_{t}=r_{t}|Z_{0:t}=z_{0:t},A_{0:t}=a_{0:t})$
$=\sum_{s_{t+1}\in S}P(S_{t+1}=s_{t+1},R_{t}=r_{t}|Z_{0:t}=z_{0:t},A_{0:t}=a_{0:t})P(Z_{t+1}=z_{t+1}|S_{t+1}=s_{t+1},Z_{0:t}=z_{0:t},A_{0:t}=a_{0:t})$
$=\sum_{s_{t+1}\in S}P(s_{t+1},r_{t}|z_{0:t},a_{0:t})P(z_{t+1}|s_{t+1})$
$=\sum_{s_{t+1}\in S}P(s_{t+1},r_{t}|~\boxed{\bigcap_{\tau\in D_{\mathcal{N}}}\{Z_{\tau}=z_{\tau},A_{\tau}=a_{\tau}\}})O_{t}(z_{t+1}|s_{t+1})$
$=P(z_{t+1},r_{t}|\ \boxed{\bigcap_{\tau\in D_{\mathcal{N}}}\{Z_{\tau}=z_{\tau},A_{\tau}=a_{\tau}\}})$
therefore, by the definition of HDP MDS, $D_{\mathcal{N}}\in\mathbf{D}_{\mathcal{H}}(h_{t})$. ∎
In fact, the MDS of POMDPs could alternatively be defined in the manner of HDPs, and vice versa. However, such definitions would complicate the analysis, at least in the sense of our work. We require the concept of HDP MDS to describe directly constructed environments: in this case, there is no need to specify hidden Markov states, as we directly construct a process where the next observation and reward depend on certain past observations and actions within the trajectory. Conversely, we need the concept of POMDP MDS to characterize the process of constructing new tasks by transforming an existing MDP environment. Here, the states provided by the original environment serve as hidden Markov states, which depend on several past observations and actions in the trajectory for memory models to predict with.
B.3 Example of an MDP Trajectory with Multiple MDS
Example B.1.
Figure 12: A deterministic MDP with all 0 reward: $S=\{0,1\},~A=\{a,b\},~\rho_{0}(0)=1$
In MDP Fig. 12 (as a special case of POMDP), given the trajectory $h_{2}=(011,bbb,00)$, then $\mathbf{D}(h_{2})=2^{\{0,1,2\}}$.
Proof.
Clearly, the union of multiple MDSs of the same trajectory remains an MDS of that trajectory. Therefore, we only need to show that $\{0\},\{1\},\{2\}\in\mathbf{D}(h_{2})$.
  • Since given $s_{2}=1$ and $1$ is an absorbing state, it is obvious that $s_{3}$ will always be $1$, $\{2\}\in\mathbf{D}(h_{2})$.
  • Given that $s_{1}=1$, then $s_{2}=1$ and $s_{3}=1$ can be determined sequentially, making $\{1\}\in\mathbf{D}(h_{2})$.
  • Moreover, since $s_{0}=0$ and $a_{0}=b$ imply $s_{1}=1$, then $\{0\}\in\mathbf{D}(h_{2})$ as well.
B.4 Explanation of Fig. 1
Figure 1: Visualization of MDSs of some example synthetic POMDP environments in Section 2. A colored block $(t,i)$ indicates that for any trajectory $h_{t}$, $i\in D^{M}\in\mathbf{D}(h_{t})$. Different colors denote distinct dependence relationships: state-only, action-only, and both.
MDP
In a general MDP, the next state-reward joint distribution depends solely on the current state and action.
T-Maze-8
T-maze-8 consists of an 8-length one-way channel with a T-junction at one end. Agents start from the end without the T-junction and receive instructions on the correct direction at the outset. To gain reward (the sole means of doing so), they must remember this direction while traversing the channel and select it accurately upon reaching the T-junction.
Under this mechanism, each state encountered while traversing the tunnel carries information from the initial instruction (information embedded in the initial state), and whether a reward is obtained at the 8th step depends on the action chosen at that final moment.
RepeatPrevious-4
RepeatPrevious-4 samples states via an iid process. Agents must repeat the state observed 4 time-steps prior. Under this mechanism, the subsequent reward hinges on both the current action and the state from 4 steps ago.
StatelessPendulum
The classical Pendulum can be modeled as an MDP, where the angular position and angular velocity define its state. The StatelessPendulum variant removes the angular position from the original Pendulum’s state representation.
As a result, agents retain a memory of all historical observations of angular velocities to reconstruct the original angular position through aggregation.
Appendix C HDP’s Transition Invariance
C.1 Independent Decision Process (IDP)
Definition C.1 (Independent Decision Process (IDP)).
An IDP $\mathcal{I}\coloneqq\langle A,\{T_{t}\}_{t=0}^{\infty}\rangle$ consists of: action set $A$, reward dynamics at time $t$: $T_{t}:A\rightarrow\Delta$.
Note: IDP is a special case of MDP with only one state, that is why $\rho_{0}$ and $S$ are omitted from the definition.
C.2 Stationary MDP
Definition C.2 (Stationary MDP).
An MDP $\mathcal{M}=\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ is stationary if $\exists T:S\times A\to\Delta_{S\times\real},~\forall t\in\mathbb{N},~T_{t}=T$.
Note: In most of the literature, the MDP is stationary by default in the form of $\langle\rho_{0},S,A,T\rangle$.
C.3 High-Order MDP
Definition C.3 (High-Order MDP).
A $k$-th ($k\in\mathbb{N}^{+}$) order MDP $\mathcal{M}_{k}\coloneqq\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ consists of: initial state distribution $\rho_{0}\in\Delta_{S}$; state set $S$; action set $A$; transition and reward dynamics at time $t$:
$T_{t}:\begin{cases}S^{k}\times A^{k}\to\Delta_{S\times\real},&(t\geq k-1)\\ S^{t+1}\times A^{t+1}\rightarrow\Delta_{S\times\real},&(t<k-1)\end{cases}$
Note: The transition dynamics of $k$-th order MDP takes states and actions from the current step back up to at most $k$ previous steps as input. It can also be defined as a special case of HDP that has a constant MDS representation $\forall h_{t},~\{t,t-1,t-2,\dots,t-k+1\}\cap\mathbb{N}\in\mathbf{D}(h_{t})$.
C.4 The Prefix Relation between Trajectories
Definition C.4 (The Prefix Relation between Trajectories).
Given an HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ with trajectory set $H$, for any $h_{t_{1}},h^{\prime}_{t_{2}}\in H,~(t_{1}\leq t_{2})$, if $h_{t_{1}}=(z_{0:t_{1}},a_{0:t_{1}},r_{0:t_{1}-1}),~h^{\prime}_{t_{2}}=(z_{0:t_{2}},a_{0:t_{2}},r_{0:t_{2}-1})$, then $h_{t_{1}}$ is a prefix of $h^{\prime}_{t_{2}}$, denoted $h_{t_{1}}\preceq h_{t_{2}}^{\prime}$.
Note: When a policy is specified, and new observations and rewards are generated step-by-step according to the transition dynamics specified by the HDP, the resulting trajectory sequence $h_{0},h_{1},h_{2},\dots$ satisfies $h_{0}\preceq h_{1}\preceq h_{2}\preceq\dots$.
C.5 Equivalence Relations Form a Lattice
In the main text, the symbol $\wedge$” is used to combine equivalence relations, and this section will provide the theoretical basis for this notation.
Definition C.5 (Lattice (via partially ordered set (poset))).
Let $\langle L,\preceq\rangle$ be a poset, where $\preceq$ denotes a partial order on the set $L$. $\langle L,\preceq\rangle$ is called a lattice if for all $a,b\in L$:
  • The supremum (denoted $a\vee b$) of $a$ and $b$ exists in $L$, which satisfies $a\preceq a\vee b,~b\preceq a\vee b$, and for any $c\in L$ with $a\preceq c$ and $b\preceq c$, it follows that $a\vee b\preceq c$.
  • The infimum (denoted $a\wedge b$) of $a$ and $b$ exists in $L$, which satisfies $a\wedge b\preceq a,~a\wedge b\preceq b$, and for any $c\in L$ with $c\preceq a$ and $c\preceq b$, it follows that $c\preceq a\wedge b$.
Definition C.6 (Lattice (via algebraic operations)).
Let $\langle L,\vee,\wedge\rangle$ be an algebraic system, where $\vee$ and $\wedge$ are binary operations on L. $\langle L,\vee,\wedge\rangle$ is called a lattice if for all $a,b,c\in L$:
  • Commutativity: $a\vee b=b\vee a$ and $a\wedge b=b\wedge a$.
  • Associativity: $(a\vee b)\vee c=a\vee(b\vee c)$ and $(a\wedge b)\wedge c=a\wedge(b\wedge c)$.
  • Absorption: $a\vee(a\wedge b)=a$ and $a\wedge(a\vee b)=a$.
Note: These operations induce a partial order $\preceq$ on $L$ via $a\preceq b\iff a\wedge b=a$ (or equivalently $a\vee b=b$), making the definitions above equivalent.
Theorem C.1 (Lattice of Equivalence Relations).
Let $S$ be a set, denote $E(S)$ the set of all equivalence relations on $S$, then $\langle E(S),\subseteq\rangle$ is a lattice.
Proof.
By the property of $\subseteq$, $\langle E(S),\subseteq\rangle$ forms a poset. For any $\phi,\psi\in E(S)$, we define:
$\phi\wedge\psi~\coloneqq\phi\cap\psi$ (1)
$\phi\vee\psi~\coloneqq\phi\cup(\psi\circ\phi)^{\ast}$ (2)
where $\forall x,y\in E(S)$, $x\circ y\coloneqq\{(a,b)|~\exists c\in S,~(a,c)\in x,~(c,b)\in y\}$ and $x^{\ast}\coloneqq x\cup x\circ x\cup x\circ x\circ x\cup\dots$.
First, we show $\phi\wedge\psi,\phi\vee\psi\in E(S)$.
  • $\phi\wedge\psi$:
    • Reflexivity: $\forall a\in S$, $(a,a)\in\phi,(a,a)\in\psi\implies(a,a)\in\phi\cap\psi=\phi\wedge\psi$.
    • Symmetry: $\forall a,b\in S,~(a,b)\in\phi\wedge\psi\implies(a,b)\in\phi,~(a,b)\in\psi\implies(b,a)\in\phi,~(b,a)\in\psi\implies(b,a)\in\phi\wedge\psi$.
    • Transitivity: $\forall a,b,c\in S,~(a,b),(b,c)\in\phi\wedge\psi\implies(a,b),(b,c)\in\phi,~(a,b),(b,c)\in\psi\implies(a,c)\in\phi,~(a,c)\in\psi\implies(a,c)\in\phi\wedge\psi$.
  • $\phi\vee\psi$:
    • Reflexivity: $\forall a\in S,~(a,a)\in\phi\implies(a,a)\in\phi\cup(\psi\circ\phi)^{\ast}=\phi\vee\psi$.
    • Symmetry: $\forall a,b\in S,~(a,b)\in\phi\vee\psi\implies(a,b)\in\phi\cup(\psi\circ\phi)^{\ast}\implies(a,b)\in\phi$ or $(a,b)\in(\psi\circ\phi)^{\ast}$.
      1. If $(a,b)\in\phi$, then $(b,a)\in\phi\implies(b,a)\in\phi\vee\psi$.
      2. If $(a,b)\in(\psi\circ\phi)^{\ast}$, then $\exists k\in\mathbb{N},~\exists x_{0:2k-1}\in S^{2k},~a=x_{0},~b=x_{2k-1},~\forall i\in\mathbb{N},i<k,~(x_{2i},x_{2i+1})\in\psi,~(x_{2i+1},x_{2i+2})\in\phi\implies(x_{2i+1},x_{2i})\in\psi,~(x_{2i+2},x_{2i+1})\in\phi\implies(b,a)\in(\psi\circ\phi)^{k}\subseteq(\psi\circ\phi)^{\ast}\subseteq\phi\vee\psi$.
    • Transitivity: $\forall a,b,c\in S,~(a,b),(b,c)\in\phi\vee\psi\implies(a,b),(b,c)\in\phi$ or $(a,b)\in\phi,~(b,c)\in(\psi\circ\phi)^{\ast}$ or $(a,b)\in(\psi\circ\phi)^{\ast},~(b,c)\in\phi$ or $(a,b),(b,c)\in(\psi\circ\phi)^{\ast}$.
      1. If $(a,b),(b,c)\in\phi$, then $(a,c)\in\phi\subseteq\phi\vee\psi$.
      2. If $(a,b)\in\phi,~(b,c)\in(\psi\circ\phi)^{\ast}$, then $(a,c)\in\phi\circ(\psi\circ\phi)^{\ast}=(\{(x,x)|x\in S\}\circ\phi)\circ(\psi\circ\phi)^{\ast}\subseteq(\psi\circ\phi)\circ(\psi\circ\phi)^{\ast}=(\psi\circ\phi)^{\ast}\subseteq\phi\vee\psi$.
      3. If $(a,b)\in(\psi\circ\phi)^{\ast},~(b,c)\in\phi$, then $\exists k\in\mathbb{N},~(a,c)\in(\psi\circ\phi)^{k}\circ\phi=(\psi\circ\phi)^{k}\subseteq(\psi\circ\phi)^{\ast}\subseteq\phi\vee\psi$.
      4. If $(a,b),(b,c)\in(\psi\circ\phi)^{\ast}$, then $(a,c)\in(\psi\circ\phi)^{\ast}\circ(\psi\circ\phi)^{\ast}=(\psi\circ\phi)^{\ast}\subseteq\phi\vee\psi$.
Then we show that such definitions of $\wedge,\vee$ meet the properties required by Def. C.5.
  • Supremum:
    • $\phi\subseteq\phi\vee\psi$:
      $\phi\subseteq\phi\cup(\psi\circ\phi)^{\ast}\implies\phi\subseteq\phi\vee\psi$.
    • $\psi\subseteq\phi\vee\psi$:
      $\psi=\psi\circ\{(x,x)|x\in S\}\subseteq\psi\circ\phi\subseteq(\psi\circ\phi)^{\ast}\subseteq\phi\cup(\psi\circ\phi)^{\ast}\implies\psi\subseteq\phi\vee\psi$.
    • $\phi\subseteq\varphi,~\psi\subseteq\varphi\implies\phi\vee\psi\subseteq\varphi$:
      $\phi\vee\psi=\phi\cup(\psi\circ\phi)^{\ast}\subseteq\varphi\cup(\varphi\circ\varphi)^{\ast}=\varphi\implies\phi\vee\psi\subseteq\varphi$.
  • Infimum:
    • $\phi\wedge\psi\subseteq\phi$:
      $\phi\wedge\psi=\phi\cap\psi\subseteq\phi$.
    • $\phi\wedge\psi\subseteq\psi$:
      $\phi\wedge\psi=\phi\cap\psi\subseteq\psi\implies\phi\wedge\psi\subseteq\psi$.
    • $\varphi\subseteq\phi,~\varphi\subseteq\psi\implies\varphi\subseteq\phi\wedge\psi$:
      $\varphi\subseteq\phi,~\varphi\subseteq\psi\implies\varphi\subseteq\phi\cap\psi=\phi\wedge\psi$.
Therefore, $\langle E(S),\subseteq\rangle$ is a lattice with $\wedge,\vee$ defined in Eq. 1, 2. ∎
Corollary C.1.
Given a trajectory set $H$ of any HDP, denote $E(H)$ the set of all equivalence relations on $H$, then $\langle E(H),\subseteq\rangle$ is a lattice.
C.6 Stationarity and Consistency
To facilitate subsequent discussions, we first present the concept of terminal trajectory.
Definition C.7 (Terminal Trajectory).
For any HDP $\mathcal{H}$ with trajectory set $H$, a trajectory $h\in H$ is a terminal trajectory if there exists no $h^{\prime}$ such that $h\preceq h^{\prime}$. The set of all terminal trajectories is denoted by $\partial H$.
Stationarity
: We use the term (intra-trajectory) stationarity to describe the invariance of the transition dynamics across steps within the same trajectory. It means that during any single run of the environment, transitions between any two steps follow the same logic. The formal definition is as follows:
Definition C.8 ($p$-Stationary).
For any HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ with trajectory set $H$, let $p:H\to p(H)$ be a map. The HDP $\mathcal{H}$ is $p$-stationary if, for all $h_{t_{1}},h_{t_{2}}\in H$ where $h_{t_{1}}\preceq h_{t_{2}}$ and $p(h_{t_{1}})=p(h_{t_{2}})$, the transition probabilities satisfy $T_{t_{1}}(z,r|z_{0:t_{1}},a_{0:t_{1}})=T_{t_{2}}(z,r|z_{0:t_{2}},a_{0:t_{2}})$ for all $z\in Z,~r\in\real$.
Note: Based on the above definitions, the following conclusions can be drawn.
  • Since $h_{t_{1}}\preceq h_{t_{2}}$, this implies there exists $h_{T}\in\partial H$ such that $h_{t_{1}}\preceq h_{t_{2}}\preceq h_{T}$. In this sense, $h_{t_{1}},h_{t_{2}}$ can be said to lie within the same trajectory $h_{T}$.
  • In this paper, the adopted $p$ satisfies $p:(z_{0:t},a_{0:t},r_{0:t-1})\mapsto(z_{t-k+1,t},a_{t-k+1,t})$ which means it considers the most recent $k$ observation-action pairs. From this, $\simeq_{k}$-invariant can be derived.
  • In this paper, we break stationarity by the following method: trajectories are divided into several stages by time intervals of equal length, with the same transition dynamics applied within each stage but potentially different ones across stages.
  • It should be noted that this is not the only way to break stationarity; we merely chose one that is relatively simple and easy to implement. The switching of transition dynamics within a trajectory can depend not only on the numerical value of the time step but also on special events occurring in the trajectory, which is quite common in real-world problems.
Consistency
We use the term (inter-trajectory) consistency to describe the invariance of the transition dynamics among different trajectories. It means that all trajectories follow the same transition logic at moments equivalent under some criterion $q$. The formal definition is as follows:
Definition C.9 ($q$-Consistent).
For any HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ with trajectory set $H$, let $q:\mathbb{N}\to q(\mathbb{N})$ be a map. The HDP $\mathcal{H}$ is $q$-consistent if, for all $h_{t_{1}},h^{\prime}_{t_{2}}\in H$ where $q(t_{1})=q(t_{2})$, the transition probabilities satisfy $T_{t_{1}}(z,r|z_{0:t_{1}},a_{0:t_{1}})=T_{t_{2}}(z,r|z^{\prime}_{0:t_{2}},a^{\prime}_{0:t_{2}})$ for all $z\in Z,~r\in\real$.
Note: Based on the above definitions, the following conclusions can be drawn.
  • Since the above definition does not restrict whether $h_{t_{1}}$ and $h^{\prime}_{t_{2}}$ form a prefix relation, this captures the meaning of ”different trajectories” in the description of the consistency concept.
  • In this paper, the adopted $q$ satisfies $q:t\mapsto\lfloor\frac{t}{n}\rfloor$ which means it treats moments within the same interval of length $n$ as equivalent. From this, $\simeq^{n}$-invariant can be derived.
  • In this paper, we break consistency using the following approach: the observation space is divided into several intervals of equal length, and the category of a trajectory is determined by the interval containing its initial observation. Trajectories within the same category share the same transition dynamics at equivalent time steps, while no such constraint applies to trajectories across different categories.
  • It should be noted that this is not the only way to break consistency: we merely chose one that is relatively simple and easy to implement. We partition trajectories into disjoint sets according to the interval containing their initial observations, which satisfy $\forall h_{t_{1}},h_{t_{2}}^{\prime}\in H,~h_{t_{1}}\sim h_{t_{2}}^{\prime}\iff\exists h\in H,~h\preceq h_{t_{1}},~h\preceq h_{t_{2}}^{\prime}$. Thus, this partitioning method enables the agent to possess sufficient information at any moment in the trajectory to determine the category of the current trajectory. However, this is not mandatory, as other ways of partitioning the trajectory set can also be adopted, in which case this property is violated.
C.7 Explanation of Tab. 1 and Fig. 3
Invariance Stationary Non-stationary
Consistent $\simeq_{k}$ $\simeq_{k}\wedge\simeq^{n}$
Non-consistent $\simeq_{k}\wedge\simeq^{\lfloor mz_{0}\rfloor}$ $\simeq_{k}\wedge\simeq^{n}\wedge\simeq^{\lfloor mz_{0}\rfloor}$
Table 1: A construction of HDPs ($S=[0,1)$) with different transition invariance properties. $k,n,m\in\mathbb{N}^{+}$, $\simeq^{\lfloor mz_{0}\rfloor}$ denotes an equivalence relation where trajectories $h,h^{\prime}$ are equivalent if their initial observations $z_{0},z_{0}^{\prime}$ lie within the same interval $\left[\frac{i}{m},\frac{i+1}{m}\right)$, $k$ is the order as shown in Fig. 2.
Figure 3: Corresponding to Tab. 1, the x-axis shows $n$-length time periods in trajectories, and the y-axis shows distinct trajectories differentiated by the $\frac{1}{m}$-length intervals that their first observations belong to. Different colors correspond to different transition and reward dynamics.
Parameter $k,n,m$
$k$ is the order of AR process, meaning $z_{t+1}$ depends on the most recent $k$ observations $z_{t-k+1:t}$.
$n$ denotes the division of the time axis into segments of length $n$, within each segment, the same (stationary) or different (non-stationary) AR coefficients can be used for transition calculations.
$m$ represents the number of intervals in the uniform partition of the observation set. A trajectory’s category is determined by the interval containing its initial observation. Across different categories, the AR coefficients used for transition calculations within the same time segment can be the same (consistent) or different (non-consistent).
Example
Given $k=2,n=1,m=2$:
Invariance Stationary Non-stationary
Consistent $z_{t+1}=w_{0}z_{t}+w_{1}z_{t-1}$ $z_{t+1}=w_{0}^{t}z_{t}+w_{1}^{t}z_{t-1}$
Non-consistent $z_{t+1}=\begin{cases}w_{0}z_{t}+w_{1}z_{t-1}~(z_{0}\leq\frac{1}{2})\\ w_{0}^{\prime}z_{t}+w_{1}^{\prime}z_{t-1}~(z_{0}>\frac{1}{2})\end{cases}$ $z_{t+1}=\begin{cases}w_{0}^{t}z_{t}+w_{1}^{t}z_{t-1}~(z_{0}\leq\frac{1}{2})\\ w_{0}^{\prime t}z_{t}+w_{1}^{\prime t}z_{t-1}~(z_{0}>\frac{1}{2})\end{cases}$
Table 2: Examples of observation generation AR processes under different stationarity and consistency conditions.
Appendix D HDP Environments Constructed from Scratch
Order Range
: $k\in\{0,1,2,3,4,5,6,7\}$.
Observation Space
: $[0,1)$.
Action Space
: $\{0,1,2,3,4,5,6,7\}$.
Trajectory Length
: $64$.
Initialization
: $z_{t}\sim U(0,1)~(0\leq t<k)$.
Coefficients
: $w_{0:7}=(8,9,10,11,12,13,14,15)/8\subset[1,2)$.
Reward
: $r_{t}=\mathbf{1}_{a_{t}=\left\lfloor 8\left(\sum_{i=0}^{k-1}w_{i}z_{t-i}\right)\mod 1\right\rfloor}~(t\geq k-1),~r_{t}=\mathbf{1}_{a_{t}=\left\lfloor 8\left(\sum_{i=0}^{t}w_{i}z_{t-i}\right)\mod 1\right\rfloor}~(t<k-1)$.
Note that $z_{t+1}=\left(\sum_{i=0}^{k-1}w_{i}z_{t-i}\right)\mod 1$ when $t\geq k-1$, otherwise sampled from $U(0,1)$.
D.1 AllEqOneLinProc(k)(k)
Coefficients
: $1_{0:7}$.
Transition
: $z_{t+1}=\left(\sum_{i=0}^{k-1}z_{t-i}\right)\mod 1$.
Description
: AR process with all coefficients set to $1$, used to test the memory model’s ability to retain observations within a finite time window of length $k$. Each observation in the window has the same weight, or in other words, decision value.
D.2 AllEqLinProc(k)(k)
Coefficients
: $w_{0:7}$.
Transition
: $z_{t+1}=\left(\sum_{i=0}^{k-1}w_{i}z_{t-i}\right)\mod 1$.
Description
: AR process with varying coefficients, used to test the performance of the memory model in consistent and stationary environments of different orders. All coefficients are no less than $1$, which is to prevent $z_{t}$ from converging to 0.
D.3 TimeEqLinProc(k)(k)
Coefficients
: Define operator $\sigma:w_{0:T}\mapsto(w_{T},w_{0:T-1})$.
Then $w_{0:7,0:7}\coloneqq(\sigma^{0}(w_{0:7}),\sigma^{1}(w_{0:7}),\sigma^{2}(w_{0:7}),\dots,\sigma^{7}(w_{0:7}))=\begin{pmatrix}8&9&10&11&12&13&14&15\\ 15&8&9&10&11&12&13&14\\ 14&15&8&9&10&11&12&13\\ 13&14&15&8&9&10&11&12\\ 12&13&14&15&8&9&10&11\\ 11&12&13&14&15&8&9&10\\ 10&11&12&13&14&15&8&9\\ 9&10&11&12&13&14&15&8\\ \end{pmatrix}/8$
Transition
: $z_{t+1}=\left(\sum_{i=0}^{k-1}w_{\lfloor t/8\rfloor,i}z_{t-i}\right)\mod 1$.
Description
: AR process with varying coefficients, used to test the performance of the memory model in consistent but non-stationary environments of different orders. Non-stationarity means that the transition functions differ between various moments within the same trajectory, i.e., the adopted AR coefficients are different. To avoid either insignificant differences in transition relationships between time segments due to an excessive number of distinct coefficients or excessive instability of the process caused by overly large coefficients, a circulant matrix construction is employed.
D.4 TrajEqLinProc(k)(k)
Coefficients
$w^{\prime}_{0:7,0:7}\coloneqq(\sigma^{0}(w_{0:7}),\sigma^{-1}(w_{0:7}),\sigma^{-2}(w_{0:7}),\dots,\sigma^{-7}(w_{0:7}))=\begin{pmatrix}8&9&10&11&12&13&14&15\\ 9&10&11&12&13&14&15&8\\ 10&11&12&13&14&15&8&9\\ 11&12&13&14&15&8&9&10\\ 12&13&14&15&8&9&10&11\\ 13&14&15&8&9&10&11&12\\ 14&15&8&9&10&11&12&13\\ 15&8&9&10&11&12&13&14\\ \end{pmatrix}/8$
Transition
: $z_{t+1}=\left(\sum_{i=0}^{k-1}w^{\prime}_{\lfloor 8z_{0}\rfloor,i}z_{t-i}\right)\mod 1$.
Description
: AR process with varying coefficients, used to test the performance of the memory models in non-consistent but stationary environments of different orders. Non-consistency means that in the same environment, an agent may face multiple possible tasks, each with its unique transition regulation; i.e., the AR coefficients used differ across sequences whose initial observations belong to different intervals. To avoid either insignificant differences in transition relationships between different trajectory categories due to an excessive number of distinct coefficients or excessive instability of the process caused by overly large coefficients, a circulant matrix construction is employed.
D.5 NoEqLinProc(k)(k)
Coefficients
: $w_{i,j,0:7}\coloneqq\sigma^{i-j}(w_{0:7}),~i,j\in\{0,1,2,3,4,5,6,7\}$.
Transition
: $z_{t+1}=\left(\sum_{i=0}^{k-1}w_{\lfloor t/8\rfloor,\lfloor 8z_{0}\rfloor,i}z_{t-i}\right)\mod 1$.
Description
: AR process with varying coefficients, used to test the performance of the memory models in non-consistent and non-stationary environments of different orders. A circulant tensor is used based on the considerations outlined in the previous subsections.
Appendix E State Aggregation
E.1 History Aggregator for State (HAS) and Optimality Preserving Property
Definition E.1 (Value Function).
Given HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ with trajectory set $H$ and a policy $\pi=\pi_{0:\pi}$, the return function is
$V^{\pi}(z_{0})\coloneqq\mathbb{E}_{h_{0:\infty}\sim(\pi_{0:\infty},T_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}$
where $h_{0,\infty}=(z_{0:\infty},a_{0:\infty},r_{0:\infty})\in\partial H$, $(\pi_{0:\infty},T_{0:\infty}|z_{0})$ represents the probability of generating a terminal trajectory that starts from $z_{0}$ through transition dynamics $T_{0,\infty}$ and policy $\pi_{0:\infty}$.
Definition E.2 (Latest Extraction Operators).
The latest state, action, and reward extraction operators at time $t$ are:
  • $\mathscr{L}_{S,t}:h_{t}\mapsto s_{t}~~(\forall t\in\mathbb{N},~\forall h_{t}\in H_{t})$;
  • $\mathscr{L}_{A,t}:h_{t+1}\mapsto a_{t}~~(\forall t\in\mathbb{N},~\forall h_{t+1}\in H_{t+1})$;
  • $\mathscr{L}_{\real,t}:h_{t+1}\mapsto r_{t}~~(\forall t\in\mathbb{N},~\forall h_{t+1}\in H_{t+1})$.
Definition E.3 (History Aggregator for State (HAS)).
For any MDP $\mathcal{M}=\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$, an HAS $\mathscr{A}_{S}\coloneqq\{\mathscr{A}_{S,t}\}_{t=0}^{\infty}$ is a series of maps in which $\mathscr{A}_{S,t}:H_{t}\to\mathscr{A}(H),~(\forall t\in\mathbb{N})$ where $\mathscr{A}(H)$ represents a set called the target state set of $\mathscr{A}_{S}$.
Note: For convenience, we assume $\mathscr{A}_{S,t}:S^{t+1}\to Z$, where $Z$ denotes the observation space of the target HDP. This means actions and rewards do not participate in constructing the current observation.
Definition E.4 (Reversibility of HAS on MDP).
An HAS $\mathscr{A}_{S}=\{\mathscr{A}_{S,t}\}_{t=0}^{\infty}$ on MDP $\mathcal{M}=\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ is reversible iff there exists a series of maps, denoted as $\mathscr{A}_{S}^{\ast}\coloneqq\{\mathscr{A}_{S,t}^{\ast}\}_{t=0}^{\infty}$, which satisfy:
$\mathscr{A}_{S,t}^{\ast}(\{\mathscr{A}_{S,\tau}(h_{\tau})\}_{\tau=0}^{t})=s_{t},~~(\forall t\in\mathbb{N})$
where $h_{\tau}\in H_{\tau},~h_{0}\prec h_{1}\prec\cdots\prec h_{t},~s_{t}=\mathscr{L}_{S,t}(h_{t})$.
The reversibility condition of HAS ensures that a decision-making algorithm can reconstruct the original MDP state from the trajectory of the HDP. This ensures that a reversible HAS does not alter the original belief state (i.e., the state of the MDP itself). Thus, the reversible HAS ensures that the inherent difficulty remains unchanged when wrapping an MDP into an HDP. Here, inherent difficulty refers to the difficulty posed to the RL algorithm itself, excluding the difficulty of extracting useful information from HDP history to construct MDP states (i.e., the difficulty for the memory model).
Definition E.5 (HAS Induced HDP).
For any MDP $\mathcal{M}=\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ with trajectory set $H=\bigsqcup_{t=0}^{\infty}H_{t}$ and an HAS $\mathscr{A}_{S}=\{\mathscr{A}_{S,t}\}_{t=0}^{\infty}$, where $\mathscr{A}_{S,t}:H_{t}\to\mathscr{A}(H)$. Applying $\mathscr{A}_{S}$ to $\mathcal{M}$ yields an HDP $\mathcal{H}=\langle\rho_{0}^{\prime},Z^{\prime},A^{\prime},\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$, where: $\rho_{0}^{\prime}=\rho_{0}\circ\mathscr{A}_{S,0}^{\ast}$, $Z=\mathscr{A}(H)$, $T^{\prime}_{t}(z_{t+1},r_{t}|z_{0:t},a_{0:t})=T_{t}(s_{t+1},r_{t}|s_{t},a_{t}),~z_{\tau}=\mathscr{A}_{S,\tau}(s_{0:\tau}),~(0\leq\tau\leq t+1)$, which is called the HDP induced from $\mathcal{M}$ by HAS $\mathscr{A}_{S}$.
Definition E.6 (HAS Induced Policy).
Given an MDP $\mathcal{M}=\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$, let $\mathcal{H}=\langle\rho_{0}^{\prime},Z^{\prime},A^{\prime},\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$ be the HDP induced by HAS $\mathscr{A}_{S}$. For any policy $\pi:S\to\Delta_{A}$ of $\mathcal{M}$, the policy $\tilde{\pi}_{t}(a_{t}|z_{0:t})\coloneqq\pi_{t}(a_{t}|s_{t})$ where $z_{\tau}=\mathscr{A}_{S,\tau}(s_{0:\tau}),~(0\leq\tau\leq t)$, is called the policy induced from $\pi$ by HAS $\mathscr{A}_{S}$.
Note: It can be stated that for any policy $\tilde{\pi}$ of HDP $\mathcal{H}$, there exists a policy $\pi$ of MDP $\mathcal{M}$ satisfying $\pi_{t}(a_{t}|s_{t})=\tilde{\pi}_{t}(a_{t}|z_{0:t}),~s_{\tau}=\mathscr{A}_{S,\tau}^{\ast}(z_{0:\tau}),~(0\leq\tau\leq t)$, such that $\tilde{\pi}$ is induced from $\pi$ by HAS $\mathscr{A}_{S}$.
Definition E.7 (MDP Optimal Policy).
For any MDP $\mathcal{M}=\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$, $\pi^{\ast}\coloneqq\pi^{\ast}_{0:\infty}$ is said to be an optimal policy of $\mathcal{M}$ if $\forall s_{0}\in S,~\pi^{\ast}\in\arg\max_{\pi}V^{\pi}(s_{0})$, where $V^{\pi}$ is the value function: $V^{\pi}(s_{0})=\mathbb{E}_{h_{0:\infty}\sim(\pi_{0:\infty},T_{0:\infty}|s_{0})}\sum_{t=0}^{\infty}r_{t}$ and $h_{0:\infty}=(s_{0:\infty},a_{0:\infty},r_{0:\infty})$.
Note: Note that the discount factor $\gamma$ can be absorbed into the time-dependent transition representation $T_{t}$ or the MDP can be restricted to a finite length before falling into an absorbing state with subsequent rewards fixed at zero. Both approaches ensure the convergence of the value function for the MDP defined above.
Definition E.8 (HDP Optimal Policy).
For any HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$, $\pi^{\ast}\coloneqq\pi_{0:\infty}$ is said to be an optimal policy of $\mathcal{H}$ if $\forall z_{0}\in Z,~\pi^{\ast}\in\arg\max_{\pi}V^{\pi}(z_{0})$, where $V^{\pi}$ is the value function $V^{\pi}(z_{0})=\mathbb{E}_{h_{0:\infty}\sim(\pi_{0:\infty},T_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}$ and $h_{0:\infty}=(z_{0:\infty},a_{0:\infty},r_{0:\infty})$.
Theorem E.1 (Optimality Preserving Property of Reversible HAS).
Given an MDP $\mathcal{M}=\langle\rho_{0},S,A,\{T_{t}\}_{t=0}^{\infty}\rangle$, let $\mathcal{H}=\langle\rho_{0}^{\prime},Z^{\prime},A^{\prime},\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$ be the HDP induced by a reversible HAS $\mathscr{A}_{S}$. For any optimal policy $\pi^{\ast}$ of $\mathcal{M}$, the policy $\tilde{\pi}^{\ast}$ induced from $\pi^{\ast}$ by $\mathscr{A}_{S}$ is an optimal policy of HDP $\mathcal{H}$.
Proof.
Since $\pi^{\ast}$ is an optimal policy of $\mathcal{M}$, then
$\pi^{\ast}\in\arg\max_{\pi}V^{\pi}(s_{0})=\arg\max_{\pi}\mathbb{E}_{h_{0:\infty}\sim(\pi_{0:\infty},T_{0:\infty}|s_{0})}\sum_{t=0}^{\infty}r_{t}$ (3)
Since $\tilde{\pi}^{\ast}$ is the policy induced from $\pi^{\ast}$ by reversible HAS $\mathscr{A}_{S}$, then
$\tilde{\pi}_{t}^{\ast}(a_{t}|z_{0:t})=\pi_{t}^{\ast}(a_{t}|s_{t})$ (4)
$z_{\tau}=\mathscr{A}_{S,\tau}(s_{0:\tau}),~s_{\tau}=\mathscr{A}_{S,\tau}^{\ast}(z_{0:\tau})$ (5)
Now the target is proving
$\tilde{\pi}^{\ast}\in\arg\max V^{\tilde{\pi}}(z_{0})=\arg\max_{\tilde{\pi}}\mathbb{E}_{h^{\prime}_{0:\infty}\sim(\tilde{\pi}_{0:\infty},T^{\prime}_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}$ (6)
Assume there is a policy $\tilde{\pi}$ such that $\exists z_{0},~V^{\tilde{\pi}}(z_{0})>V^{\tilde{\pi}^{\ast}}(z_{0})$, which is
$\mathbb{E}_{h^{\prime}_{0:\infty}\sim(\tilde{\pi}_{0:\infty},T^{\prime}_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}>\mathbb{E}_{h^{\prime}_{0:\infty}\sim(\tilde{\pi}^{\ast}_{0:\infty},T^{\prime}_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}$ (7)
By Eq. 5, there is a bijection $\alpha$ between $H$ (trajectory set of $\mathcal{M}$) and $H^{\prime}$ (trajectory set of $\mathcal{H}$):
$\alpha:(s_{0:t},a_{0:t},r_{0:t})\mapsto(\mathscr{A}_{S,t}(s_{0:t}),a_{0:t},r_{0:t}),~\alpha^{-1}:(z_{0:t},a_{0:t},r_{0:t})\mapsto(\mathscr{A}_{S,t}^{\ast}(z_{0:t}),a_{0:t},r_{0:t})$ (8)
Denote the MDP policy from which $\tilde{\pi}$ is induced as $\pi$, we can rewrite Eq. 7:
$\mathbb{E}_{h_{0:\infty}\sim(\pi_{0:\infty},T_{0:\infty}|s_{0})}\sum_{t=0}^{\infty}r_{t}>\mathbb{E}_{h_{0:\infty}\sim(\pi^{\ast}_{0:\infty},T_{0:\infty}|s_{0})}\sum_{t=0}^{\infty}r_{t}$ (9)
Therefore $\exists s_{0}\in S,~V^{\pi}(s_{0})>V^{\pi^{\ast}}(s_{0})$, contradicts Eq. 3.
Thus, assumption Eq. 7 can not be true, which means $\forall z_{0}\in Z,~\tilde{\pi}^{\ast}\in\arg\max_{\tilde{\pi}}V^{\tilde{\pi}}(z_{0})$, $\tilde{\pi}^{\ast}$ is an optimal policy of HDP $\mathcal{H}$. ∎
Note: The proof of this property is not provided in the HAS paper (Wang and Li 2025).
E.2 Convolution Based Reversible HAS
General Form
$z_{t}=\mathscr{A}_{S,t}(s_{0:t})\coloneqq w_{0:t}\ast s_{0:t}=\sum_{i=0}^{t}w_{i}s_{t-i}$ (10)
$\begin{pmatrix}z_{0}\\ z_{1}\\ z_{2}\\ \dots\\ z_{t}\end{pmatrix}=\begin{pmatrix}w_{0}&0&0&\dots&0\\ w_{1}&w_{0}&0&\dots&0\\ w_{2}&w_{1}&w_{0}&\dots&0\\ \dots&\dots&\dots&\dots&\dots\\ w_{t}&w_{t-1}&w_{t-2}&\dots&w_{0}\\ \end{pmatrix}\cdot\begin{pmatrix}s_{0}\\ s_{1}\\ s_{2}\\ \dots\\ s_{t}\end{pmatrix}$ (11)
$z_{0:t}=W_{t}s_{0:t}$ (12)
Therefore, when $w_{0}\neq 0$,
$s_{0:t}=W_{t}^{-1}z_{0:t}$ (13)
Thus, using the concept of POMDP MDS, $W_{t}^{-1}$ determines all observations $z_{t}$ required to decode the current state $s_{t}$.
Example
Given $w_{0:\infty}=(1,w,0^{\ast})$, then
$W_{4}=\begin{pmatrix}1&0&0&0\\ w&1&0&0\\ 0&w&1&0\\ 0&0&w&1\end{pmatrix};~W_{4}^{-1}=\begin{pmatrix}1&0&0&0\\ (-w)^{1}&1&0&0\\ (-w)^{2}&(-w)^{1}&1&0\\ (-w)^{3}&(-w)^{2}&(-w)^{1}&1\end{pmatrix}$ (14)
This example serves as the basis for the visualization in Fig. 6. When $w$ is positive, the coefficients of each observation alternate in sign; when $w$ is negative, all coefficients of the observations are positive.
Figure 6: Visualization of MDSs of convolution. Left figure coefficients: $1,w$; Right figure coefficients: 1,w1,-w, $(w>0)$. In $W^{-1}_{t}$, positive elements are red, negative blue, and darker colors indicate larger absolute values of the coefficient.
E.3 Generalized Convolution Implementation
In fact, besides using the addition and scalar multiplication of vectors in Euclidean space, there are other ways to implement these two operations of convolution, as long as they still form a linear space. For example, given any Abelian group operator $S\times S\to S$, a simple way is to pick an arbitrary bijection $\eta:S\to S$ and define the addition operator to be $\oplus:(s,s^{\prime})\mapsto\eta^{-1}(\eta(s)+\eta(s^{\prime}))$, the scalar multiplication operator to be $\odot:(w,s)\mapsto\eta^{-1}(w\cdot\eta(s))$. It suffices to verify that such a definition still satisfies the requirements of a linear space.
Discrete State Space
If the state space is discrete, there are two approaches to aggregate states for constructing an HDP: either continuousizing the states and processing them in the same way as for continuous state spaces, or preserving the discreteness of the observation space.
The continuousization method is simple to implement. It merely requires mapping discrete states to integer codes and then treating them as real numbers.
Preserving discreteness, however, does not allow such an operation. In this case, it is necessary to define addition and scalar multiplication operations on discrete sets (e.g., $\{0,1,2,\dots,N-1\}$). In this case, modular-N addition and modular-$N$ scalar multiplication can be used to replace the corresponding operations in the continuous state space, and the invertibility condition of $W_{t}$ becomes that $w_{0}$ is coprime with $N$. Since all structures of Abelian groups on finite sets are known, they are isomorphic to the direct sum of certain cyclic groups. Even non-Abelian groups have complete classification theorems. Coupled with the arbitrary choice of bijections on the state set, this provides a great variety of options for constructing convolution-based HAS.
Appendix F Reward Redistribution
F.1 Reward Redistribution and Optimality Preserving Property
Definition F.1 (Return-Equivalent HDP).
Two HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ and $\mathcal{H}^{\prime}=\langle\rho_{0}^{\prime},Z^{\prime},A^{\prime},\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$ are return-equivalent if $Z=Z^{\prime},~\rho_{0}=\rho_{0}^{\prime},~A=A^{\prime},~\sum_{r\in\real}T_{t}(z,r|z_{0:t},a_{0:t})=\sum_{r\in\real}T_{t}^{\prime}(z,r|z_{0:t},a_{0:t})$ and for all $z_{0}\in Z$ and any policy $\pi=\pi_{0,\infty}$, $V_{\mathcal{H}}^{\pi}(z_{0})=V^{\pi}_{\mathcal{H}^{\prime}}(z_{0})$.
Note: For two return-equivalent HDPs, their observation transition dynamics are identical, while their reward mechanisms are not entirely the same: although the rewards of each trajectory have different temporal distributions, their total amounts remain the same.
Definition F.2 (Reward Redistribution).
A reward redistribution is a map $\mathscr{R}:\mathcal{H}\mapsto\mathcal{H}^{\prime}$ where $\mathcal{H},\mathcal{H}^{\prime}$ are return-equivalent HDPs.
Theorem F.1 (Optimality Preserving Property of Reward Redistribution).
For any reward redistribution $\mathscr{R}$ and any HDP $\mathcal{H}$, $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle,~\mathcal{H}^{\prime}=\mathscr{R}(\mathcal{H})=\langle\rho_{0}^{\prime},Z^{\prime},A^{\prime},\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$ have the same set of optimal policies.
Proof.
For any optimal policy $\pi$ of $\mathcal{H}$, $\forall z_{0}\in Z$,
$\pi^{\ast}\in\arg\max_{\pi}V^{\pi}_{\mathcal{H}}(z_{0})=\arg\max_{\pi}\mathbb{E}_{h_{0:\infty}\sim(\pi_{0:\infty},T_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}$ (15)
Since $\mathcal{H},\mathcal{H}^{\prime}$ are return-equivalent,
$\mathbb{E}_{h_{0:\infty}\sim(\pi_{0:\infty},T_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}=\mathbb{E}_{h^{\prime}_{0:\infty}\sim(\pi_{0:\infty},T^{\prime}_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}^{\prime}$ (16)
where $h_{0:\infty}=(z_{0:\infty},a_{0:\infty},r_{0\infty}),~h^{\prime}_{0:\infty}=(z_{0:\infty},a_{0:\infty},r^{\prime}_{0\infty})$. Therefore,
$\pi^{\ast}\in\arg\max_{\pi}\mathbb{E}_{h^{\prime}_{0:\infty}\sim(\pi_{0:\infty},T^{\prime}_{0:\infty}|z_{0})}\sum_{t=0}^{\infty}r_{t}^{\prime}=\arg\max_{\pi}V^{\ast}_{\mathcal{H}^{\prime}}(z_{0})$ (17)
$\pi^{\ast}$ is an optimal policy of $\mathcal{H}^{\prime}$.
By symmetry of return-equivalence relation of $\mathcal{H},\mathcal{H}^{\prime}$, for any optimal policy $\tilde{\pi}^{\ast}$ of $\mathcal{H}^{\prime}$, $\tilde{\pi}^{\ast}$ is also an optimal policy of $\mathcal{H}$. ∎
F.2 Reward Delay
Definition F.3 (Reward Delay).
Reward delay $\mathscr{D}_{k}$ is a reward redistribution that maps HDP $\mathcal{H}=\langle\rho_{0},Z,A,\{T_{t}\}_{t=0}^{\infty}\rangle$ (with trajectory set $H$) to $\mathcal{H}^{\prime}=\langle\rho_{0},Z,A,\{T_{t}^{\prime}\}_{t=0}^{\infty}\rangle$ (with trajectory set $H^{\prime}$) such that for any trajectory $h_{0:t}=(z_{0:t},a_{0:t},r_{0:t})\in H$, its corresponding trajectory $h^{\prime}_{0:t}=(z_{0:t},a_{0:t},r^{\prime}_{0:t})$ in $H^{\prime}$ satisfies (if $0\leq T\leq\infty$ is the max timestep)
$r^{\prime}_{0:k-1}=0,~r^{\prime}_{k:T-2}=r_{0:T-k-2},~r^{\prime}_{T-1}=\sum_{i=0}^{k-1}r_{T-i-1}$
The transition dynamics $T^{\prime}_{t}$ satisfy
$T_{t}^{\prime}(z_{t+1},0|z_{0:t},a_{0:t})=\sum_{r_{t}\in\real}T_{t}(z_{t+1},r_{t}|z_{0:t},a_{0:t}),~(t<k)$
$T_{t}^{\prime}(z_{t+1},r^{\prime}_{t}|z_{0:t},a_{0:t})=\mathbf{1}_{r_{t-k}=r^{\prime}_{t}}\sum_{r_{t}\in\real}T_{t}(z_{t+1},r_{t}|z_{0:t},a_{0:t}),~(k\leq t\leq T-2)$
$T_{T-1}^{\prime}(z_{T},r^{\prime}_{T-1}|z_{0:T-1},a_{0:T-1})=\mathbf{1}_{\sum_{i=0}^{k-1}r_{T-i-1}=r_{T-1}^{\prime}}\sum_{r_{T-1}\in\real}T_{T-1}(z_{T},r_{T-1}|z_{0:T-1},a_{0:T-1})$
Note: It should be noted that delaying all rewards of each time step to appear after a certain number of moments is not the only way to implement delayed rewards. Rewards can also be delayed in a certain proportion, where both the proportion and the delay time can be variable.
If there is a discount factor $\gamma$, the delayed reward should be
$r^{\prime}_{0:k-1}=0,~r^{\prime}_{k:T-2}=\frac{r_{0:T-k-2}}{\gamma^{k}},~r^{\prime}_{T-1}=\sum_{i=0}^{k-1}\frac{r_{T-1-i}}{\gamma^{i}}$
It is obvious that reward delay $\mathscr{D}_{k}$ is a reward redistribution.
Appendix G HDP Environments Derived from Existing
G.1 StateConv(k)(k)
Coefficients
: $w_{0}^{k}=1,~w_{!}^{k}=1-\frac{1}{2^{k}},~(k\leq 4),~w_{1}^{5}=1$.
Transition
: $z_{t}=w_{0}^{k}s_{t}+w_{1}^{k}s_{t-1},~(t>0)$, $z_{0}=w_{0}^{k}s_{0}$.
Description
: Corresponding to Section E.2, StateConv(k)(k) is a series of environment wrappers. By wrapping a given MDP environment, it constructs the observation sequence of the HDP by convolving the states of the MDP, while keeping the reward mechanism unchanged. As $k$ increases, the current observation contains more information about historical states, which requires the memory model to remember more historical state information to decode the current Markov state, thus making the problem more difficult for the memory model. This wrapper can be applied to various tasks with continuous observation spaces, provided that the states support scalar multiplication and addition operations.
G.2 RewardDelay(k)(k)
Transition
: The reward appearing at moment $t$ in the given environment is delayed to moment $t+k$, with the rewards at times 0 to k1k-1 set to zero. Meanwhile, the rewards in the $k$ moments before termination are all moved to the termination time. The delay is compensated according to the power of the discount.
Description
: Corresponding to Section F.2, RewardDelay(k)(k) delays rewards while ensuring that the optimal policy remains unchanged, still being a Markov policy in the original MDP. In this case, for agents using a memory model, the difficulty does not lie in remembering historical observations, constructing belief states, and making better decisions, but rather in learning a policy that relies solely on the current observation. This series of wrappers tests the ability of the memory model to ignore irrelevant information during the learning process. As $k$ increases, the number of steps by which rewards are delayed increases; since rewards are no longer generated immediately, the amount of irrelevant information that the memory needs to ignore also grows, which increases the fitting difficulty for the memory model.
Appendix H Experiments Settings and Further Discussion
H.1 Settings
CPU
: Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz; 28 Cores, 112 Threads.
GPU
: NVIDIA GeForce RTX 3080, Driver Version: 550.127.05, CUDA Version: 12.4; 8 GPUs.
Memory
: 1TiB.
OS
: 24.04.1 LTS (Noble Numbat).
Packages
: python[3.10.0]; numpy[1.25.0]; torch[2.7.1+cu126]; ray[2.46.0]; gymasium[1.0.0]; scipy[1.15.3]; popym.
H.2 Configs and Hyperparameters
Repeated Runs
: $10$ times for each experiment.
Memory Models
: The standard baseline implementations from the POPGym library were used.
Algorithm
: PPO using ray-rllib, modified from the implementation in POPGym.
Discount Factor $\gamma$
: 0 for LinProc series and RewardWhenInside wrapped with StateConv series (Since maximizing the reward at each individual step in these environments ensures the maximization of the total reward over the entire trajectory, $\gamma$ is set to 0), $0.9$ for other environments.
RLlib Hyperparameters
: Provided in the code. Mainly in train scripts: run_xxx.py
H.3 Further Discussion
Impact of Non-Consistency and Non-Stationarity
(a) Consistent, Stationary
(b) Consistent, Non-stationary
(c) Non-consistent, Stationary
(d) Non-consistent, Non-stationary
Figure 9: Linear processes with different coefficients and different transition invariance.
As can be observed in Fig. 9, the impact of non-stationarity on the performance of the memory model is not significant. We hypothesize that there may be the following reasons: 1. The overall differences in the transition mechanisms used in different time segments are not obvious enough. To keep the transition process relatively stable, the coefficients we adopted are in the range of $1$ to $2$, which makes the impact of non-stationarity on performance insufficiently significant. 2. In our code implementation, we have adopted sufficiently large learnable parameters (of preprocessing module, postprocessing module, and memory model itself), which enables various memory models to handle such a degree of non-stationarity over these AR sequences.
We have also analyzed the reasons why non-consistency affects performance and provided the following possible explanations: The way non-consistency is implemented in this paper is to determine the category of the current trajectory through the first observation, and then select different transition mechanisms among different categories. Essentially, this method introduces a dependency on the initial moment for each step. This makes the problem’s MDS more complex, implying that the memory model needs to keep track of the initial observation at all times, which poses a severe challenge to its performance.
As for the scenario where both non-consistency and non-stationarity are satisfied, the performance degradation of the memory model is actually less significant than in the environment with only non-consistency. We conjecture that this may result from the fact that we only adopt a limited number of coefficients and use a cyclic mechanism to construct the transition relationships across different time segments of trajectories belonging to different categories. This implies that the boundaries of transition mechanisms between different categories are blurred, allowing a segment of the transition mechanism from certain category trajectories to appear in other time periods of trajectories from other categories. Consequently, the memory model can achieve better generalization across these trajectories of different categories.
What Kind of MDP Environment is Suitable for Wrapping
(a) RewardWhenInside_p
(b) RewardWhenInside_n
(c) NormalizedCartPole_p
(d) NormalizedCartPole_n
Figure 10: StateConv series with increasing coefficient $w_{1}$.
For the two environment series StateConv(k)(k) and RewardDelay(k)(k), in addition to wrapping RewardWhenInside and NormalizedCartPole, we also wrapped other environments such as Acrobot, MountainCar and did experiments on them. When applying these wrappers with increasing difficulty levels to these environments, we cannot guarantee that the plotted performance curves will be as perfectly monotonically decreasing as those on wrapped RewardWhenInside. Instead, non-ideal situations may occur, such as those shown in the bottom two figures of Fig. 10, where the performance of the memory model may improve despite an increase in difficulty level at certain stages.
We would like to emphasize that this phenomenon is not counterintuitive, but rather highly reasonable. It reflects the deep-seated intrinsic nature of the state representations and transition dynamics of the wrapped problems. Taking StateConv(k)(k) as an example, the ideal scenario is that all optimal policies rely on recovering the original MDP state representation from the observation sequence. This holds naturally in the wrapped RewardWhenInside environment. However, for other environments—especially those MDP environments whose state representations contain a large amount of redundant information—decision-making in these wrapped MDP environments does not necessarily require perfect recovery of the original state. In fact, the optimal policy can even be well-represented using only observations. This situation runs counter to our original intention: by obtaining observation sequences through convolution, we aim to force the memory model to learn to decode states from observation sequences, thereby increasing the difficulty of the problem.
Thus, if we intend to use the wrapper series to generate a set of tasks with increasing difficulty, we need to carefully examine the characteristics and intrinsic structures of the MDPs to be wrapped. We should select MDPs with sufficiently compact state representations and relatively independent distributions of states across steps as the objects of wrapping. Only in this way can the optimal policy representation learned by the memory model depend on the perfect recovery of the original MDP states.